40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9078     Entry time: Wed Aug 28 03:28:00 2013
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Playing with DRMI, some ASS automation prep 

[Rana, Jenne]

We did lots of poking around with the DRMI tonight.  I should elog more in the morning, but the most important points are:

Locking settings same as elog 9068, except PRCL gain changed to 0.035, and the FMs that are triggered.  PRCL tonight had FM2,3,6 triggered.  MICH had FM1,2,3,7 triggered.  SRCL had FM1,2 triggered.  Engaging the MICH boosts helped make things more quiet, so that some of the SRC boosts could be enabled.  Still not as good of lock stretches as Koji got last Friday (elog 9060). 

REFL55 and REFL11 were still saturating (only during acquisition), after the optical path changes I did last week (elog 9043).  We reduced the REFL55 whitening slider from 15 dB to 6 dB (but forgot to compensate with digital gain), to keep the counts (as seen on DTT time series, binning off) to less than ~20,000 counts.  REFL11 is still saturating, and we're not sure why, since it's slider gain is 0 dB.  To be investigated.

I was prepping the ASS to be more conveniently put into a wrapper script, which could be called from the IFO Configure screen.  This involved adding PRCL to the burt .req and .snap files, as well as modifying the scripts a little bit to include PRCL as an option.  I ended up changing the script names from DITHER_Arm_ON.py and DITHER_Arm_OFF.py to DITHER_ASS_ON.py adn DITHER_ASS_OFF.py, since they are no longer restricted to being arms-only.  You must still provide an argument to the script, to tell it which degree of freedom you want to activate.  I also changed the save offsets scripts.  The way they were, the X and Y arms just had separate hard-coded scripts, with no convenient way to incorporate PRCL.  I merged them (including PRCL) into WRITE_ASS_OFFSETS.py, which you must now provide the DoF as an argument.  I tested these new scripts on all 3 of the DoFs, and made changes to the ASS screen, so it now calls only the new scripts.  It should now be easy to incorporate future ASS modifications.

Rana was in the middle of modifying the ASS model to include SRCL, and we also need to include MICH.  The ASS model is not compile-ready, so don't compile it!!  If you need to compile the ASS, please save what's there as a different name, and do an "svn up" to get the latest working version.


We suspected that there might be angular drive issues with the SRM (it was wiggling a lot). We checked the damping via step responses - all Qs were less than 10. Then we found that the INPUT button on the SRM PIT OL was OFF (why ???). After turning this back on it behaved better. We measured the loop shape and found that the UGF was 7 Hz; good. Need to work on some loop shaping for this guy. Its just 1/f out to 300 Hz right now. UGF should be made a little lower so that we can stably turn on the Bounce/Roll notches and a ~50 Hz low pass filter.

Most importantly, the F2A filters need to be measured and implemented. They are a few years old.

ELOG V3.1.3-